# Hungarian Research Network (HUN-REN) Laboratory on Engineering and Management Intelligence Research Group of Operations Research and Decision Systems

### v1

@@ Abstract
The organisers of major sports competitions use different policies with respect to constraints in the group draw. Our paper aims to rationalise these choices by analysing the trade-off between attractiveness (the number of games played by teams from the same geographic zone) and fairness (the departure of the draw mechanism from a uniform distribution). A parametric optimisation model is formulated and applied to the 2022 FIFA World Cup draw. A flaw of the draw procedure is identified: the pre-assignment of the host to a group implies additional but unnecessary distortions. All Pareto efficient sets of draw constraints are determined via simulations. The proposed framework can be used to find the optimal draw rules of a tournament and justify the distortion of the draw procedure for the stakeholders.
@@
Keywords: OR in sports; constrained assignment; Monte Carlo simulation; tournament design; uniform draw
@@ 1 Introduction
Several sports tournaments start with a group stage: the teams are allocated into some groups to play a round-robin contest and select the best teams in the group that qualify for the next stage. The process of assigning the teams to the groups is called the group draw. The group draws of major sports competitions are among the most followed randomisation procedures in practice. For example, the video of the 2022 FIFA World Cup draw (https://www.youtube.com/watch?v=PgV_z-TA0aw)—the draw that will serve as the case study for our optimisation model—has been watched 4.7 million times.
The most important requirements towards a group draw are balancedness, transparency, and uniform distribution (Guyon,, 2015). Balance means that the groups must be roughly at the same competitive level; the probability of qualification is not allowed to depend substantially on the outcome of the draw. Transparency is connected to the implementation of the draw; it requires drawing not too many balls from a small number of urns, carried out by celebrities and broadcast live across media channels. Uniform distribution is a fairness principle: every possible draw outcome should have the same probability to occur.
@@
However, the set of teams playing in the same group may be subject to further restrictions. In particular, geographic separation is an important issue in the Olympic Games and World Championships because the stakeholders want to see unique matches between teams coming from different geographic regions. If geographical constraints need to be met, the three basic criteria may be violated since no mechanism is known that can satisfy all of them under arbitrary restrictions. The Fédération Internationale de Football Association (FIFA), the international governing body of (association) football, has often sacrificed balancedness in the history of the FIFA World Cup (Jones,, 1990; Rathgeber and Rathgeber,, 2007; Guyon,, 2015). After Guyon, (2015) has identified several flaws in the 2014 FIFA World Cup draw, FIFA has adopted the so-called Skip mechanism for both the 2018 and 2022 FIFA World Cup draws (Guyon,, 2018; Csató, 2023a, ).
@@
  • The model is applied to the 2022 FIFA World Cup draw, which involves extensive numerical simulations for 32 reasonable sets of draw restrictions (Section 4);
@@
  • Depending on the measure of non-uniformity, only three or four sets of constraints could be optimal according to the preferences of the organiser (Section 4.3).
Thus, we have a clear policy recommendation to improve the 2022 FIFA World Cup draw, as well as the draw of several major sports tournaments. The suggested approach can also be directly used to optimise the set of draw constraints before the draw takes place, and justify the unavoidable distortion of the transparent draw procedure for all stakeholders.
@@ 2 Related literature
As we have seen in Section 1, balancedness, transparency, and uniform distribution can be easily guaranteed in any group draw in the absence of geographical separation. Hence, the interesting case is when geographic restrictions apply. Since transparency cannot be sacrificed for the sake of other goals in practice, FIFA has struggled with balancedness and uniform distribution in the history of the FIFA World Cup draw. The draw of the 1990 (Jones,, 1990), 2006 (Rathgeber and Rathgeber,, 2007), and 2014 (Guyon,, 2015) events have been revealed to violate uniform distribution. The 2014 FIFA World Cup draw has been quite imbalanced, too (Guyon,, 2015). Therefore, Guyon, (2014)—the full preliminary version of the published paper Guyon, (2015)—has proposed three fairer and still transparent draw procedures, which inspired FIFA to adopt the Skip mechanism for the 2018 and 2022 FIFA World Cup draws (Guyon,, 2018). Tijms, (2015) underlines the importance of transparency by discussing the credibility of accusations that the draw for the quarterfinals in the 2012/13 UEFA Champions League has been manipulated.
Lapré and Palazzolo, (2023) empirically assess competitive imbalance between the FIFA World Cup groups from 1954 to 2022 to show its serious consequences with respect to the chance of reaching the quarterfinals. Consequently, reducing imbalance should be an important aim for the organiser. Cea et al., (2020) develop a mixed integer linear programming model to improve the balance between the groups with retaining geographical separation. Laliena and López, (2019) propose two draw systems for allocating 32 teams into eight balanced groups subject to arbitrary constraints. The two mechanisms represent the trade-off between balancedness and transparency. Csató, 2023a uncovers that the official seeding policy of the 2022 FIFA World Cup draw has increased imbalance between the groups; a better alternative would have been assigning the placeholders according to the highest-ranked potential winner rather than automatically seeding them in the weakest pot (see Section 3.2). The methods of Laliena and López, (2025) for balancing the strengths of groups containing three teams start from the ideas of Laliena and López, (2019) but contain exogenous lower and upper bounds. The authors also examine non-uniformity but do not consider changing the geographic criteria.
@@
However, none of these works discuss the possible gains from slacking the geographic constraints, despite that they have never been used in handball and volleyball. According to our knowledge, only one result exists on this issue in the literature. The scope for improving fairness distortions in the UEFA Champions League Round of 16 draw (Klößner and Becker,, 2013) is found to be limited through a better draw procedure but a quantitatively large reduction can be achieved by allowing one match (instead of zero) between two teams from the same country (Boczoń and Wilson,, 2023, Section 6.2). Nonetheless, there are fundamental differences: (1) the fairness distortion measure of Boczoń and Wilson, (2023) does not quantify the level of non-uniformity (it can exceed zero even if the draw procedure has uniform distribution); (2) the setting of a bipartite graph with 16 nodes (UEFA Champions League Round of 16 draw) is much simpler than the eight groups of four teams in the 2022 FIFA World Cup draw; and (3) the UEFA Champions League Round of 16 draw has used a distinct mechanism (Csató, 2025d, ).
@@
The current paper is related to the numerous papers investigating various design issues in the FIFA World Cup, too. Monks and Husch, (2009) estimate the effect of home continent and seeding on the probability of qualification. Scarf and Yusof, (2011) explore the impact of the seeding policy without considering geographical constraints. Csató, 2023b shows that the probability of qualifying for the FIFA World Cup is not independent of the confederation to which a particular team is assigned; for example, a South American team could have tripled its chances by playing in Asia. Guyon, (2020), Chater et al., (2021), and Stronka, (2024) examine the connection between the schedule of group matches and their competitiveness. Stronka, (2020) and Csató, 2025b focus on the temptation to deliberately lose a match in the last round of the group stage. Krumer and Moreno-Ternero, (2023) and Csató et al., (2025) suggest fair allocations of qualifying berths among the continents. Guyon, (2020), Guajardo and Krumer, (2024), and Csató and Gyimesi, (2025) propose alternative formats for the 2026 FIFA World Cup, the first event with 48 teams. A recent survey of these studies can be found in Devriesere et al., (2025).
@@ 3.1 Draw procedures
Draw constraints such as geographic separation make an assignment of the teams to the groups invalid if it violates at least one constraint. The statistical challenge resides in simulating (closely) from the uniform distribution with a sequential and entertaining method. Governing bodies in sports use two procedures for this purpose, the Drop mechanism and the Skip mechanism (Csató, 2025d, ). The Skip mechanism is more popular as it is employed in the FIBA Basketball World Cup, the FIFA World Cup, and the UEFA Nation League, among other tournaments. Its underlying idea is straightforward: if a draw condition occurs or is anticipated to occur, the computer used to assist with the draw indicates the next available group in alphabetical order for the drawn team (UEFA,, 2024).
@@
Consider the draw presented in Example 1. There are four possible outcomes if the group labels are ignored but one of them (when teams 2, 4, 6 are assigned to the same group) is invalid, resulting in three feasible assignments. In particular, team 1 can be in the same group with (a) teams 3 and 6; (b) teams 4 and 5; or (c) teams 4 and 6. Thus, teams 1 and 4 play in a group with a probability of 2/3 in a uniform draw.
@@
3.2 The rules of the 2022 FIFA World Cup
The 2022 FIFA World Cup draw took place on 1 April 2022. The 32 national teams were divided into four pots of eight teams each. Pot 1 contained the host Qatar, which was automatically assigned to Group A, as well as the seven highest-ranked teams in the FIFA World Rankings of March 2022. Pot 2 consisted of the next eight highest-ranked teams, and Pot 3 consisted of the next eight. Finally, Pot 4 contained the five lowest-ranked nations (ignoring the host Qatar) and three placeholders, the two winners of the inter-confederation play-offs and one winner of a UEFA play-off still to be played. One team was assigned from each pot to a group to ensure roughly the same competitive level in all groups.
FIFA has six confederations: AFC (Asian Football Confederation), CAF (Confederation of African Football), CONCACAF (Confederation of North, Central American and Caribbean Association Football), CONMEBOL (Confederación Sudamericana de Fútbol), OFC (Oceania Football Confederation), and UEFA (Union of European Football Associations). In the qualification for the FIFA World Cup, each confederation received a predetermined number of slots, possible including shared slots allocated by inter-confederation play-offs. In order to guarantee geographic separation and maximise the number of inter-continental games in the group stage, the following restrictions were used (FIFA,, 2022):
@@ 3.2 The rules of the 2022 FIFA World Cup
The placeholders for the inter-confederation play-offs were considered for all confederations involved. The matches of the qualification played after the draw are shown in Table 1.
Table 2: The geographical composition of the pots in the 2022 FIFA World Cup draw
@@
  • Placeholders for the inter-confederation play-offs are taken for both related confederations into account, see Table 1.
Table 2 shows the number of teams in the four pots from the six confederations. Note that at most one OFC team may play in the group stage, hence, there is no need to avoid a match between two OFC teams by an additional restriction.
@@ 3.3 The optimisation model
Table 3: Sets of possible draw constraints in the 2022 FIFA World Cup draw
@@
The 2022 FIFA World Cup draw used the Skip mechanism, starting with Pot 1 and ending with Pot 4. According to Section 3.1, this procedure is non-uniform that might threaten the fairness of the draw. In the absence of a uniform and transparent draw procedure (Roberts and Rosenthal,, 2024), a reasonable step towards uniformity is offered by ignoring some of Constraints A–E. In particular, Constraints A to E can be separately imposed or disregarded,implying 32 different sets of draw constraints $\mathcal{C}$. They are listed in Table 3 together with their notation by integers from 0 to 31.
@@
Therefore, FIFA has two reasonable goals when designing the rules of the draw:
@@
The non-uniformity of the draw under a set of constraints $\mathcal{C}$ can be quantified in several ways; we will use two measures. The first is the mean absolute bias of all nonzero matching probabilities, measured in percentage points:
@@
The second is the maximum absolute bias of all matching probabilities, measured in percentage points:
@@
where parameter $\alpha$ reflects the priority of the decision-maker with respect to the objectives. In other words, the aim is minimising the weighted average of non-uniformity and intra-confederation games. Multiplier 10 has been added to $\Delta(\mathcal{C})$ in (3) based on the case study of the 2022 FIFA World Cup draw as the values of $10\Delta(\mathcal{C})$ and $\Psi(\mathcal{C})$ are roughly of the same magnitude, but does not affect the set of Pareto efficient solutions.

### v2

# How to optimise tournament draws: The case of the FIFA World Cup

@@ Abstract
The organisers of major sports competitions use different policies with respect to constraints in the group draw. Our paper aims to rationalise these choices by analysing the trade-off between attractiveness (the number of games played by teams from the same geographic zone) and fairness (the departure of the draw mechanism from a uniform distribution). A parametric optimisation model is formulated and applied to the 2018 and 2022 FIFA World Cup draws. A flaw of the draw procedure is identified: the pre-assignment of the host to a group unnecessarily increases the distortions. All Pareto efficient sets of draw constraints are determined via simulations. The proposed framework can be used to find the optimal draw rules and justify the non-uniformity of the draw procedure for the stakeholders.
@@
Keywords: constrained assignment; Monte Carlo simulation; OR in sports; tournament design; uniform draw
@@ 1 Introduction
Several sports tournaments start with a group stage: the teams are allocated into some groups to play a round-robin contest—where each team plays the same number of games against every other team—and select the best teams in the group that qualify for the next stage. The process of assigning the teams to the groups is called the group draw. The group draws of major sports competitions are among the most followed randomisation procedures in practice. For example, the video of the 2022 FIFA World Cup draw (https://www.youtube.com/watch?v=PgV_z-TA0aw)—the draw that will serve as the case study for our optimisation model—has been watched 4.9 million times as of 30 September 2025.
According to recent papers in operational research (see Section 2) and the policy of governing bodies in sports, the most important requirements towards a group draw are probably balancedness, transparency, and uniform distribution (Guyon,, 2015). Balance means that the groups must be roughly at the same competitive level; the probability of qualification is not allowed to depend substantially on the outcome of the draw. Transparency is connected to the implementation of the draw; it requires drawing not too many balls from a small number of urns, carried out by celebrities and broadcast live across media channels. Uniform distribution is a fairness principle: every possible draw outcome should have the same probability to occur.
@@
However, the set of teams playing in the same group may be subject to further restrictions. In particular, geographic separation is an important issue in the Olympic Games and World Championships because the stakeholders want to see unique matches between teams coming from different geographic regions. If geographical constraints need to be met, the three basic criteria may be violated since no mechanism is known that can satisfy all of them under arbitrary restrictions. The Fédération Internationale de Football Association (FIFA), the international governing body of (association) football, has often sacrificed balancedness in the history of the FIFA World Cup (Jones,, 1990; Rathgeber and Rathgeber,, 2007; Guyon,, 2015). After Guyon, (2015) has identified several flaws in the 2014 FIFA World Cup draw, FIFA has adopted the so-called Skip mechanism for both the 2018 and 2022 FIFA World Cup draws (Guyon,, 2018; Csató,, 2023).
@@
  • The model is applied to the 2018 and 2022 FIFA World Cup draws, which involve extensive numerical simulations for 32 reasonable sets of draw restrictions (Section 4);
@@
  • Depending on the composition of the draw pots and the measure of non-uniformity, at most five sets of constraints could be optimal according to the preferences of the organiser (Section 4.3).
Thus, we have a clear policy recommendation to improve the FIFA World Cup draw, as well as the draw of several major sports tournaments. The suggested approach can also be directly used to optimise the set of draw constraints before the draw takes place, and justify the unavoidable distortion of the transparent draw procedure for all stakeholders.
@@ 2 Related literature
As we have seen in Section 1, balancedness, transparency, and uniform distribution can be easily guaranteed in any group draw in the absence of geographical separation. Hence, the interesting case is when geographic restrictions apply. Since transparency cannot be sacrificed for the sake of other goals in practice, FIFA has struggled with balancedness and uniform distribution in the history of the FIFA World Cup draw. The draws of the 1990 (Jones,, 1990), 2006 (Rathgeber and Rathgeber,, 2007), and 2014 (Guyon,, 2015) events were revealed to violate uniform distribution. The 2014 FIFA World Cup draw was quite imbalanced, too (Guyon,, 2015). Therefore, Guyon, (2014)—the full preliminary version of the published paper Guyon, (2015)—has proposed three fairer and still transparent draw procedures, which inspired FIFA to adopt the Skip mechanism for the 2018 and 2022 FIFA World Cup draws (Guyon,, 2018). Tijms, (2015) underlines the importance of transparency by discussing the credibility of accusations that the draw for the quarterfinals in the 2012/13 UEFA Champions League was manipulated.
Lapré and Palazzolo, (2023) empirically assess competitive imbalance between the FIFA World Cup groups from 1954 to 2022 to show its serious consequences with respect to the chance of reaching the quarterfinals. Consequently, reducing imbalance should be an important aim for the organiser. Cea et al., (2020) develop a mixed integer linear programming model to improve the balance between the groups with retaining geographical separation. Laliena and López, (2019) propose two draw systems for allocating 32 teams into eight balanced groups subject to arbitrary constraints. The two mechanisms represent the trade-off between balancedness and transparency. Csató, (2023) uncovers that the official seeding policy of the 2022 FIFA World Cup draw has increased imbalance between the groups; a better alternative would have been assigning the placeholders according to the highest-ranked potential winner rather than automatically seeding them in the weakest pot (see Section 3.2). The methods of Laliena and López, (2025) for balancing the strengths of groups containing three teams start from the ideas of Laliena and López, (2019) but contain exogenous lower and upper bounds. The authors also examine non-uniformity but do not consider changing the geographic criteria.
@@
However, none of these works discuss the possible gains from slacking the geographic constraints, despite that they have never been used in handball and volleyball. According to our knowledge, only one result exists on this issue in the literature. The scope for improving fairness distortions in the UEFA Champions League Round of 16 draw (Klößner and Becker,, 2013) is found to be limited through a better draw procedure but a quantitatively large reduction can be achieved by allowing one match (instead of zero) between two teams from the same country (Boczoń and Wilson,, 2023, Section 6.2). Nonetheless, there are fundamental differences: (1) the fairness distortion measure of Boczoń and Wilson, (2023) does not quantify the level of non-uniformity (it can exceed zero even if the draw procedure has a uniform distribution over all valid assignments); (2) the setting of a bipartite graph with 16 nodes (UEFA Champions League Round of 16 draw) is much simpler than the eight groups of four teams in the 2022 FIFA World Cup draw; and (3) the UEFA Champions League Round of 16 draw has used a distinct mechanism (Csató, 2025c, ).
The new incomplete round-robin format of UEFA club competitions also calls for a draw. Since the 2024/25 season, the UEFA Champions League starts with a league phase, where 36 teams play against eight opponents each and are ranked in a single table. Devriesere et al., 2025b present an integer program to check whether the schedule can be completed if a match is drawn. Guyon et al., (2025) investigate four different draw procedures for the league phase. They are implemented using integer programming to express the draw constraints. The authors compare these mechanisms by Monte Carlo simulations to evaluate their fairness and compute the matchup probabilities—what we will do for the 2018 and 2022 FIFA World Cups under 32 reasonable sets of draw constraints.
@@
The current paper is related to the numerous papers investigating various design issues in the FIFA World Cup, too. Monks and Husch, (2009) estimate the effect of home continent and seeding on the probability of qualification. Scarf and Yusof, (2011) explore the impact of the seeding policy without considering geographical constraints. Guyon, (2020), Chater et al., (2021), and Stronka, (2024) examine the connection between the schedule of group matches and their competitiveness. Stronka, (2020) focus on the temptation to deliberately lose a match in the last round of the group stage. Krumer and Moreno-Ternero, (2023) suggest fair allocations of qualifying berths among the continents. Guyon, (2020) and Guajardo and Krumer, (2024) propose alternative formats for the 2026 FIFA World Cup, the first event with 48 teams. A recent survey of these studies can be found in Devriesere et al., 2025a .
Finally, we develop an optimisation problem that allows the organiser to maximise its objective function. This is a standard approach in operational research applied to sports problems. Ribeiro et al., (2023) construct tournament schedules to minimise the number of breaks, the number of rounds where breaks occur, the difference between the number of breaks for the teams, and the carryover effect. Csató et al., (2024) compare 12 valid schedules for the UEFA Champions League group stage with respect to the probability of weakly (where one team is indifferent) and strongly (where both teams are indifferent) stakeless matches. Van Bulck et al., (2026) propose a heuristic bilevel optimization method to ensure a balanced emphasis on physical abilities and cognitive skills in rogaining, an orienteering running sport. Daltrophe et al., (2025) study the complexity of manipulating knockout tournaments for various types of manipulation, such as in favour of a single player achieving the highest score, or in favour of a given set of players achieving a higher score than the complementary set.
@@ 3.1 Draw procedures
Draw constraints such as geographic separation make an assignment of the teams to the groups invalid if it violates at least one constraint. The statistical challenge resides in simulating (closely) from the uniform distribution with a sequential and entertaining method. Governing bodies in sports use two procedures for this purpose (Csató, 2025c, ). Among them, the Skip mechanism is more popular as it is employed in the FIBA Basketball World Cup, the FIFA World Cup, and the UEFA Nations League, among other tournaments. Its underlying idea is straightforward: if the violation of a draw constraint occurs or is anticipated to occur, the computer used to assist with the draw indicates the next available group in alphabetical order for the drawn team (UEFA,, 2024).
@@
Consider the draw presented in Example 1. There are four possible outcomes if the group labels are ignored, but one of them (when teams 2, 4, 6 are assigned to the same group) is invalid, resulting in three feasible assignments. In particular, team 1 can be in the same group with (a) teams 3 and 6; (b) teams 4 and 5; or (c) teams 4 and 6. Thus, teams 1 and 4 play in a group with a probability of 2/3 in a uniform draw.
@@
3.2 The rules of the 2018 and 2022 FIFA World Cup draws
After Guyon, (2015) identified a number of flaws in the 2014 FIFA World Cup draw, the French mathematician Julien Guyon managed to convince FIFA to reform the draw of the FIFA World Cup (Guyon,, 2018). Therefore, first in the history, the draw pots were constructed by considering the historical performances of the national teams, while geographic separation was achieved by draw constraints in both the 2018 and 2022 tournaments.
In particular, the 32 national teams were divided into four pots of eight teams each. Pot 1 contained the host, automatically assigned to Group A, as well as the seven highest-ranked teams in the FIFA World Rankings. Pot 2 consisted of the next eight highest-ranked teams, and Pot 3 consisted of the next eight. Finally, Pot 4 contained the eight lowest-ranked nations in 2018, and the five lowest-ranked nations and three placeholders—the two winners of the inter-confederation play-offs and one winner of a UEFA play-off still to be played—in 2022. One team was assigned from each pot to a group to ensure roughly the same competitive level in all groups.
FIFA has six confederations: AFC (Asian Football Confederation), CAF (Confederation of African Football), CONCACAF (Confederation of North, Central American and Caribbean Association Football), CONMEBOL (Confederación Sudamericana de Fútbol), OFC (Oceania Football Confederation), and UEFA (Union of European Football Associations). In the qualification for the FIFA World Cup, each confederation received a predetermined number of slots, possibly including shared slots allocated by inter-confederation play-offs. In order to guarantee geographic separation and maximise the number of inter-continental games in the group stage, the following restrictions were used (FIFA,, 2017, 2022):
@@ 3.2 The rules of the 2018 and 2022 FIFA World Cup draws
In the 2022 FIFA World Cup draw, the placeholders for the inter-confederation play-offs were considered for all confederations involved. The qualification games played after the draw are shown in Table 1.
Table 2: The geographical composition of the pots in the recent FIFA World Cups
a 2018 FIFA World Cup draw
Confederation Pot 1 Pot 2 Pot 3 Pot 4 Sum
AFC 1 0 1 4 5
CAF 0 0 3 2 5
CONCACAF 0 1 1 1 3
CONMEBOL 2 3 0 0 5
UEFA 6 4 3 1 14
b 2022 FIFA World Cup draw
@@
  • Placeholders for the inter-confederation play-offs are taken for both confederations into account, see Table 1.
Table 2 shows the number of teams in the four pots from the six confederations. In 2018 (Table 2.a), no OFC team qualified for the FIFA World Cup. In 2022 (Table 2.b), at most one OFC team could have played in the group stage, hence, there was no need to avoid a match between two OFC teams by an additional restriction.
@@ 3.3 The optimisation model
Table 3: Sets of possible draw constraints in the recent FIFA World Cup draws
@@
Both the 2018 and 2022 FIFA World Cup draws used the Skip mechanism, starting with Pot 1 and ending with Pot 4. According to Section 3.1, this procedure is non-uniform, which might threaten the fairness of the draw. In the absence of a uniform and transparent draw procedure (Roberts and Rosenthal,, 2024), a reasonable step towards uniformity is offered by ignoring some of Constraints A–E. In particular, Constraints A to E can be separately imposed or disregarded, implying 32 different sets of draw constraints $\mathcal{C}$. They are listed in Table 3 together with their notation by integers from 0 to 31.
@@
Therefore, we assume that FIFA has two reasonable goals when designing the rules of the draw:
@@
The non-uniformity of the draw under a set of constraints $\mathcal{C}$ can be quantified in several ways; we will use two measures. The first is the mean absolute bias of all nonzero matchup probabilities, measured in percentage points:
@@
The second is the maximum absolute bias of all matchup probabilities, measured in percentage points:
@@
where parameter $\alpha$ reflects the priority of the decision-maker with respect to the objectives. In other words, the aim is minimising the weighted average of non-uniformity and intra-confederation games. Multiplier 10 is added to $\Delta(\mathcal{C})$ in (3) based on our two case studies as the values of $10\Delta(\mathcal{C})$ and $\Psi(\mathcal{C})$ are roughly of the same magnitude.
We will determine the set of Pareto efficient draw constraints that can be the optimal solutions to optimisation problems (3) and (4) for a given value of parameter $\alpha$. Naturally, they might be different for the 2018 and 2022 FIFA World Cups, but the multiplier 10 in (3) does not affect them. The organiser could explain only these choices to the public if the decision criteria are the non-uniformity of the draw and the number of intra-confederation group games. It is straightforward to obtain the optimal solution for any exogenous value of $\alpha$ because 32 sets of draw constraints are considered in each case; thus, the objective function should be evaluated only in these 32 elements of set $\mathcal{C}$.